### v2

# Rethinking the Sampling Criteria in Reinforcement Learning for LLM Reasoning: A Competence-Difficulty Alignment Perspective

Abstract
Reinforcement learning exhibits potential in enhancing the reasoning abilities of large language models, yet it is hard to scale for the low sample efficiency during the rollout phase. Existing methods attempt to improve efficiency by scheduling problems based on problem difficulties. However, these approaches suffer from unstable and biased estimations of problem difficulty and fail to capture the alignment between model competence and problem difficulty in RL training, leading to suboptimal results. To tackle these limitations, this paper introduces Competence-Difficulty Alignment Sampling (CDAS), which enables accurate and stable estimation of problem difficulties by aggregating historical performance discrepancies of problems. Then the model competence is quantified to adaptively select problems whose difficulty is in alignment with the model’s current competence using a fixed-point system. Experimental results across a range of challenging mathematical benchmarks show that CDAS achieves great improvements in both accuracy and efficiency. CDAS attains the highest average accuracy against baselines and exhibits significant speed advantages compared to Dynamic Sampling, a competitive strategy in DAPO, which is 2.33 times slower than CDAS.
Figure 1: Average accuracy (left) and training GPU hours (right) of different sampling strategies.
1 Introduction
Advanced large language models (LLMs) exemplified by DeepSeek-R1 guo2025deepseek and OpenAI O1 openaio1 demonstrate remarkable performance in challenging tasks like mathematics. As a core technology in their reports, the Reinforcement Learning algorithm (RL), such as Proximal Policy Optimization schulman2017ppo (PPO) and Group Relative Policy Optimization shao2024deepseekmathgrpo (GRPO), is employed to amplify the reasoning capabilities of the models. It works by utilizing a verifier as the reward model to guide the generation of high-quality reasoning chains without the need for data annotation. Despite the promise, the RL training is costly and hard to scale, particularly due to its low sample efficiency during the rollout phase. Recent studies yu2025dapo ; zeng2025simplerl ; bae2025online indicate that sampling overly difficult problems often results in no correct chains, while sampling overly simple problems contributes little to model capabilities, leading to computational waste. Consequently, a host of efforts are devoted to exploring sampling strategies for more efficient and stable RL training.
Existing strategies draw inspiration from Curriculum Learning (CL) bengio2009curriculum ; narvekar2020curriculum , scheduling data based on problem difficulty to enhance training stability and efficiency. Curriculum Sampling Strategy team2025kimi relies on prior difficulty labels, which are excessively offline, neglecting the inherent capabilities of the model. Dynamic Sampling used by DAPOyu2025dapo demonstrates promising results by oversampling and filtering out problems with the pass rate equal to 1 and 0, which incurs substantial rollout overhead and compromises the training efficiency. Prioritized Sampling Strategy used by Kimi k1.5team2025kimi records the latest pass rate of each problem during training and adaptively assigns a higher sampling probability to those with lower pass rates. Overall, the pass rate has been widely adopted as a proxy for modeling problem difficulty.
Figure 2: The variation in pass rate of problems in palin GRPO on Qwen-2.5 7Byang2024qwen2p5 using MATH dataset.
However, these strategies tend to be suboptimal due to two main issues: (1) Unstable and Biased Estimations of Problem Difficulty Using Single-Step Pass Rate. As our experiment of training Qwen-2.5 7Byang2024qwen2p5 with MATH datasethendrycksmathdataset2021 shows (fig. 2), the pass rate of individual problems exhibits considerable fluctuations throughout the training process, leading to unstable estimations of problem difficulty, consistent with several prior workszheng2023secrets ; peng2023stabilizing . Moreover, focusing solely on the final pass rate can introduce a difficulty bias. For instance, the pass rate of problem 2 at step 5 happens to surpass that of problem 1, despite their distinct trajectories and difficulty levels. This demonstrates that the pass rate at a single step fails to capture the true complexity and learning dynamics associated with each problem. (2) Failing to Appropriately Capture the Alignment between Model Competence and Problem Difficulty. A common strategy, such as in curriculum sampling and prioritized sampling, is to assign higher sampling probabilities to more difficult problems with lower pass rates. A practical consequence is that overemphasizing difficult samples often leads to the selection of many zero-gradient problems with a pass rate of 0 in GRPO training, limiting the training efficiency. However, a more appropriate approach, as advocated in the curriculum learningbengio2009curriculum ; narvekar2020curriculum , is to prioritize problems that are more aligned with the current level of competency of the model for a more efficient and effective training.
To address the issues above, we propose Competence-Difficulty Alignment Sampling (CDAS) for RL training, dynamically sampling problems whose difficulty matches the model competence at the step level. The core intuition behind CDAS is twofold: (1) instead of relying solely on the pass rate at a single step, an accumulative estimation that incorporates all historical information tends to yield a more stable assessment of problem difficulty; and (2) explicitly modeling model competence to measure its alignment with problem difficulty, thereby enabling more effective sampling decisions. Specifically, we model problem difficulty as the trajectory of performance discrepancy over training steps, where each point reflects the gap between the expected and actual pass rate. Then we use the centroid of this trajectory to provide a stable and accurate assessment of problem difficulty. Further, model competence is defined as the negative expected difficulty across all problems, and the absolute difference between model competence and a problem’s difficulty is used to quantify their alignment. Considering the dynamics of RL training, the competence-difficulty alignment estimation above is further formulated as a difficulty-based fixed-point system, which can iteratively converge and ensures the stability of training with theoretical guarantees.
To valid the effectiveness of CDAS, we conduct GRPO training on Qwen2.5-7Byang2024qwen2p5 using a series of sampling strategies team2025kimi ; yu2025dapo . Results across 7 comprehensive mathematical reasoning benchmarkshendrycksmathdataset2021 ; lewkowycz2022solvingminerva ; he2024olympiadbench ; cobbe2021traininggsm8k , show that CDAS consistently outperforms powerful baselines (see fig. 1), achieving the highest average accuracy of $46.77\%$. Compared to Dynamic Sampling, a highly competitive baseline, CDAS achieves slightly better performance $+0.12\%$ while considerably reducing the training step time overhead by $57.06\%$. Additionally, CDAS continually improves performance on more challenging benchmarks, achieving an accuracy of $11.77\%$ on AIME25, which surpasses the plain GRPO by $+4.79\%$ and Dynamic Sampling by $+2.19\%$. Subsequent analysis demonstrates that CDAS flexibly allocates computational resources by successfully sampling valuable problems that exhibit a pass rate not equal to 0 or 1. Further experiments on code generation tasks and larger 14B models also highlight the generalization ability of our method. The contributions of this paper can be summarized as follows:
  • We identify and analyze the limitations of existing sampling strategies from a new perspective, highlighting the importance of stable difficulty estimation and dynamic competence-difficulty alignment in RL training for LLMs.
  • We introduced Competence-Difficulty Alignment Sampling (CDAS), adaptively selecting problems that match the model competence, which is grounded in a theoretically guaranteed fixed-point system. Extensive experiments validate its effectiveness and efficiency.
2 Related Work
RL for LLMs reasoning. Reinforcement learning (RL) has been widely adopted to enhance the reasoning abilities of LLMs, especially in mathematics and programming tasks he2024olympiadbench ; trinh2024solving ; jain2025livecodebench ; penedo2025codeforces . Actor-critic-based methods, such as Proximal Policy Optimization (PPO) schulman2017ppo , utilize a value model to estimate the value function, guiding the policy updates. On the other hand, REINFORCE-based methods rely on policy gradients without a value model. Group Relative Policy Optimization (GRPO) shao2024deepseekmathgrpo normalizes rewards within a group of generated outputs, eliminating the need for a separate value model. REINFORCE++ hu2025reinforce++ enhances the classical REINFORCE algorithm williams1992simple by incorporating optimization techniques from PPO. Dynamic Sampling Policy Optimization (DAPO) yu2025dapo introduces several optimizations to enhance training efficiency and stability in long-CoT reasoning tasks.
Sampling Strategies for RL training. Effective sampling strategies are crucial for efficient RL training with LLMs. Coarse-grained curriculum learning team2025kimi ; xie2025logic gradually increases problem difficulty based on predefined labels. LIMR li2025limr introduces Learning Impact Measurement (LIM) to select problems that align with the model’s learning trajectory. Prioritized Sampling team2025kimi tracks the pass rate for each problem and samples problems in proportion to their failure rates. This approach directs the model’s focus toward more challenging problems. Dynamic Sampling yu2025dapo continues to sample problems within a batch until their pass rates are neither 0 nor 1. While this ensures that the model focuses on problems that are neither too easy nor too hard, it can lead to extremely heavy computational overhead as the training steps increase.
3 Preliminary
Group Relative Policy Optimization GRPO utilizes group-based advantage without a value model, thereby reducing computational overhead. Formally, given a problem $x$, the correct answer $y$, and a group of sampled responses $\{\hat{y_{i}}\}_{i=1}^{G}$ with their corresponding rewards $\{r_{i}\}_{i=1}^{G}$, GRPO calculates the advantage by normalizing the rewards within each group. The original GRPO objective employs a sample-level loss calculation which potentially introduces length biases yu2025dapo ; liu2025understanding ; chu2025gpg , so we utilize a token-level policy gradient loss as our objective function:
$\mathcal{J}_{\textrm{GRPO}}(\theta)=\mathbb{E}_{[x\sim\mathcal{D},\{\hat{y_{i}}\}_{i=1}^{G}\sim\pi_{\textrm{old}}(\cdot|x)]}$ (1)
$\frac{1}{\sum_{i=1}^{G}|\hat{y_{i}}|}\sum_{i=1}^{G}\sum_{t=1}^{|\hat{y_{i}}|}\bigg{(}\min\Big{(}r_{i,t}(\theta)\hat{A}_{i,t},\text{clip}\Big{(}r_{i,t}(\theta),1-\epsilon,1+\epsilon\Big{)}\hat{A}_{i,t}\Big{)}-\beta\mathbb{D}_{\text{KL}}[\pi_{\theta}\parallel\pi_{\text{ref}}]\bigg{)},$
where
$r_{i,t}(\theta)=\frac{\pi_{\theta}(\hat{y}_{i,t}\mid x,\hat{y}_{i,<t})}{\pi_{\theta_{\text{old}}}(\hat{y}_{i,t}\mid x,\hat{y}_{i,<t})},\quad\hat{A}_{i,t}=\frac{r_{i}-\text{mean}(\{r_{i}\}_{i=1}^{G})}{\text{std}(\{r_{i}\}_{i=1}^{G})}.$ (2)
Rule-Based Reward The use of reward model usually leads to reward hacking problem gao2023scaling ; weng2024rewardhack , so we use a rule-based reward function. The reward is computed using the following rule:
$r(y,\hat{y})=\begin{cases}1&\mathrm{is\_equivalent}(y,\hat{y})\\ 0&\mathrm{otherwise}\end{cases}$ (3)
Notably, we do not employ a format reward in GRPO implementation. Prior research indicates that strict format constraints may limit the upper bound of model performance zeng2025simplerl ; singh2023beyond ; wang2024planning . Therefore, we directly use the final accuracy as the reward.
4 Competence-Difficulty Alignment Sampling
In this section, we will systematically elaborate on the framework of Competence-Difficulty Alignment Sampling (CDAS), including modeling the problem difficulty from the perspective of trajectory, measuring the alignment between model and problems, and how to incorporate our CDAS into iterative GRPO training as a fixed-point system.
4.1 Modeling Problem Difficulty as a Trajectory
Figure 3: Pass Rate vs Step.
We start with a simple example in fig. 3, where the pass rate of a problem increases from 0 to 1 as the training step increases. Measuring difficulty based solely on the pass rate fails to provide a stable estimation due to the significant fluctuations in its pass rate along with the inherent limitations of information at a single step. Intuitively, by taking into account the historical performance across all previous samplings, we can obtain a much more stable and informative estimation. We plot the average historical pass rate up to the current step in fig. 3, which results in a much smoother curve that reflects the underlying historical trend.
Inspired by this, we turn to model the difficulty $D$ of a problem $x$ using the trajectory of the model’s performance discrepancy on the problem. Specifically, for the model $M_{n}$ at the $n$-th training step, its performance discrepancy $d$ on a $x$ is defined as
$d_{n}(x)=\hat{P}_{M_{n}}(y|x)-P_{M_{n}}(y|x),$ (4)
where $P_{M_{n}}(y|x)$ represents the actual probability of the model solving the problem $x$, estimated by the pass rate $s(x)$. $\hat{P}_{M_{n}}(y|x)$ represents the expected probability that the model can solve $x$. Then, given the existing trajectory $\{d_{1}(x),d_{2}(x),...,d_{n}(x)\}$, we quantify the difficulty $D$ of the problem using the centroid of the trajectory
$D_{n}(x)=\frac{1}{n}\sum_{k=1}^{n}\left(\hat{P}_{M_{k}}(y|x)-P_{M_{k}}(y|x)\right).$ (5)
The centroid considers the cumulative performance discrepancies over multiple training steps, providing a more accurate and robust measure of problem difficulty. The concept of performance discrepancy can be seen as a derivation from the $1-s(x)$ approach in Prioritized Samplingteam2025kimi , serving as a refined estimation of difficulty at each individual training step. The $1-s(x)$ metric can be regarded as a special case of the performance discrepancy, where the model is assumed to be capable of solving all problems. The introduction of $\hat{P}_{M_{n}}(y|x)$ amplifies the influence of steps where the actual pass rate deviates significantly from the model’s expectation.
The estimation of $\hat{P}_{M_{n}}(y|x)$ is supposed to account for both the problem difficulty and the model competence. Here we employ the sigmoid function, which is widely used in probability modeling, to describe such probability
$\hat{P}_{M_{n}}(y|x)=\frac{1}{1+e^{-(C_{n-1}-D_{n-1}(x))}}=\sigma(C_{n-1}-D_{n-1}(x)),$ (6)
where $C_{i-1}$ represents the model’s competence at the previous step, quantified as the negative expectation of the difficulties over all problems
$C_{n-1}=-\mathbb{E}_{x}[D_{n-1}(x)].$ (7)
When the model competence exceeds the problem difficulty, the expected probability of the problem being solved is higher. Otherwise, it is lower when the competence is lower than the difficulty.
4.2 Step-Level Adaptive Sampling based on Quantified Alignment
From eq. 6, we can see that $C_{n-1}-D_{n-1}(x)$ quantifies the degree of deviation between the model’s competence and the problem difficulty. As a result, we naturally introduce the definition of the alignment between a single problem $x$ and the model $M_{n}$ as
$\mathcal{A}(x,M_{n})=|C_{n-1}-D_{n-1}(x)|.$ (8)
The smaller the value of $\mathcal{A}$, the better the problem difficulty matches the current competence of the model. At each training step, we adopt a symmetric sampling method to construct the problem set $B=B^{-}\cup B^{+}$ from problems with the smallest $\mathcal{A}$ for rollout, where
$B^{-}$ $=\operatorname{argmin}_{|B|/2}\,\mathcal{A}(x,M_{n}),\quad\text{s.t.}\quad C_{n-1}-D_{n-1}(x)\leq 0,$ (9)
$B^{+}$ $=\operatorname{argmin}_{|B|/2}\,\mathcal{A}(x,M_{n}),\quad\text{s.t.}\quad C_{n-1}-D_{n-1}(x)>0.$ (10)
Compared to directly selecting the problems with the smallest $\mathcal{A}$, such symmetric method maintains a balanced proportion of hard and easy problems in $B$, promoting training stability.
4.3 Reinforcement Learning with CDAS
Algorithm 1 Competence-Difficulty Alignment Sampling in GRPO
1:Training set $\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{n},y_{n})\}$, Model $M_{0}$, total steps $K$, batch size $|B|$
2:Initialize $C_{0}=0$, $t_{j}=0$, $D_{t_{j}}(x_{j})=0\,(j=1,2,\ldots,N)$
3:for $n=1$ to $K$ do
4:     for $j=1$ to $N$ do
5:         $\mathcal{A}(x_{j},M_{n-1})\leftarrow|C_{n-1}-D_{t_{j}}(x)|.$
6:     end for
7:     Sample $B=B^{-}\cup B^{+}$ based on $\mathcal{A}$
8:     for $(x_{j},y_{j})$ in $B$ do  // GRPO Rollout
9:         Compute pass rate $s_{n}(x_{j})$
10:         Compute rewards $r$ for each sampled response
11:         $t_{j}\leftarrow t_{j}+1$
12:         $D_{t_{j}}(x_{j})\leftarrow\frac{t_{j}-1}{t_{j}}\cdot D_{t_{j}-1}(x_{j})+\frac{1}{t_{j}}\cdot(\sigma(C_{n-1}-D_{t_{j}-1}(x_{j}))-s_{n}(x_{j}))$
13:     end for
14:     $C_{n}\leftarrow-\mathbb{E}_{x}[D_{t_{j}}(x)]$
15:     $M_{n+1}\leftarrow$ Update the policy model $M_{i}$ by maximizing the GRPO objective
16:end for
Practically, for each training step, we first measure the alignment between the previous difficulty and model competence, and then obtain the problem set for rollout based on the symmetric sampling above. After one step, the model’s competence and the difficulty of the problems are dynamically updated. The whole process repeats until the training converges, illustrated in algorithm 1. Note that the update of problem difficulty can be viewed as a fixed-point system, formulated as
$\begin{cases}D^{*}(x)=\sigma(C^{*}-D^{*}(x))-S^{*}(x),\ \ x\in\mathbb{X}\\ C^{*}=-\mathbb{E}_{x}[D^{*}(x)]\end{cases}$ (11)
Due to the properties of the sigmoid function which is a contraction mapping, the system will converge to a unique solution. Refer to appendix C for a detailed proof of the convergence.
Note that since $|B|$ is usually much smaller than the size of the training set, performing a full update of problem difficulties at each step will lead to heavy computational overhead. Instead, for problem $x_{j}$, we record the number of times it is sampled as $t_{j}$ and update its difficulty only when it is sampled. Specifically, at the $i$-th step, the problem difficulty can be updated iteratively by
$D_{t_{j}}(x_{j})=\frac{t_{j}-1}{t_{j}}\cdot D_{t_{j}-1}(x_{j})+\frac{1}{t_{j}}\cdot(\sigma(C_{n-1}-D_{t_{j}-1}(x_{j}))-s_{n}(x_{j})).$ (12)
5 Experiments
5.1 Setups
Dataset Following zeng2025simplerl , we utilize the MATH dataset hendrycksmathdataset2021 for RL training, including 7,500 training samples and 4,500 test samples. We reserve its MATH500 subset as the validation set for RL training.
Baselines We compare CDAS against a range of powerful baselines: (1) Random Sampling, which directly samples the problem randomly from the training dataset and can be viewed as plain GRPO training. (2) Curriculum Sampling team2025kimi , which uniformly samples easy problems first, then turns to harder ones. Here, we use the difficulty level tags included in the MATH dataset, ranging from $1$ to $5$, and conduct training from the middle checkpoint in Random Sampling on problems with difficulty level $\geq 4$. (3) Prioritized Sampling, which is adopted in Kimi k1.5team2025kimi that tracks the pass rate $s$ of each problem and then samples problems based on $1-s$.(4) Dynamic Sampling, adopted in DAPOyu2025dapo , which over-samples and filters out problems with the pass rate equal to $1$ or 00.
Training Details We conduct GRPO training experiments on Qwen-2.5-7B yang2024qwen2p5 using the veRL framework sheng2024hybridflowverl . Following the training recipe in zeng2025simplerl , we set the batch size $|B|=1024$, generating $8$ rollouts for each problem with temperature $1.0$ and maximum response tokens $4096$. The training runs for a total $110$ steps, which equals training $10$ epochs on the whole training set. Considering the stability of iterations, we adopt a warm-up strategy in CDAS that ensures problems are sampled randomly within the first $11$ steps to obtain more accurate initial values for problem difficulties. Refer to appendix A for more training details.
Evaluation For a comprehensive evaluation, we select $7$ mathematical reasoning tasks, including AIME 24&25, AMC23, MATH500 hendrycksmathdataset2021 , Minerva Math lewkowycz2022solvingminerva , Olympiad Bench he2024olympiadbench and GSM8K cobbe2021traininggsm8k . Since the number of problems in AIME24&25 and AMC23 is relatively small, we report the Avg@32 metric to reduce randomness. Standard accuracy is reported for the remaining tasks. More details are shown in appendix B.
Table 1: Performance comparison across different sampling methods on various math benchmarks. Metrics are Avg@32 for AIME/AMC and standard accuracy for others. We present the best results in bold and the second with underline.
Steps Method Average Step Time (s) AIME24 (Avg@32) AIME25 (Avg@32) AMC23 (Avg@32) MATH 500 Minerva Math Olympiad Bench GSM8K Avg.
0 Qwen2.5-7B-base - 0.06 0 9.92 47 17.65 14.52 81.5 24.38
55 Random Sampling 1236 10.00 7.29 49.92 74.20 36.76 39.26 91.66 44.73
Curriculum Sampling 1236 10.00 7.29 49.92 74.20 36.76 39.26 91.66 44.73
Prioritized Sampling 1509 11.77 7.08 49.92 74.00 39.34 40.15 92.04 44.90
Dynamic Sampling 2875 12.19 7.92 51.88 75.00 38.97 38.52 91.36 45.12
CDAS (Ours) 1420 11.56 7.71 49.92 76.20 38.97 40.44 92.19 45.28
110 Random Sampling 1245 12.29 6.98 53.98 75.2 37.13 40.00 92.95 45.50
Curriculum Sampling 1375 12.71 7.19 53.98 76.0 38.60 39.56 92.57 45.80
Prioritized Sampling 1670 15.10 9.27 54.61 75.0 37.50 39.56 91.51 46.08
Dynamic Sampling 3745 15.10 9.58 53.44 77.2 39.34 39.56 92.34 46.65
CDAS (Ours) 1608 14.90 11.77 52.03 75.4 40.44 40.89 91.96 46.77
5.2 Main Results
The main results are summarized in table 1, with training curves illustrated in LABEL:fig:train. Besides the final checkpoint, we also report the intermediate checkpoint at the $55$-th step. The accuracy curve on each benchmark can be found in appendix D. From the results, we have the following findings:
CDAS outperforms plain GRPO and other baselines, achieving the highest average accuracy of $45.28\%$ at the 55-th step and $46.77\%$ the 110-th step. We also find that the improvements offered by Curriculum Sampling ($+0.30\%$ at the 110-th step) and Prioritized Sampling ($+0.58\%$ at the 110-th step) over the Random Sampling baseline are limited. It indicates that relying solely on prior difficulty labels or single-step pass rates tends to be suboptimal.
CDAS demonstrates a substantial efficiency advantage over Dynamic Sampling. While Dynamic Sampling achieves performance comparable to CDAS ($46.65\%$ vs. $46.77\%$) by oversampling and excluding problems with pass rates of 1 or 0, it incurs significantly greater computational overhead, requiring $\mathbf{2.33}$X more computation than CDAS. This result clearly highlights the superior efficiency of CDAS to Dynamic Sampling.
CDAS consistently enhances the performance on more challenging benchmarks. On benchmarks like AIME24/25, Minerva Math and Olympiad Bench, CDAS demonstrates remarkable improvements. For instance, CDAS achieves $11.77\%$ accuracy on AIME25 and $40.89\%$ on Olympiad Bench, notably higher than baselines.
The reward curve in CDAS initially grows and then converges to a median value. From LABEL:fig:train, after 30 steps, the amplitude of reward narrows to within the range of $0.4$ to $0.6$ and then converges around $0.5$. This indicates that CDAS effectively selects problems with difficulties appropriate for the model. On the other hand, since Prioritized Sampling tends to select harder problems, its reward demonstrates a continuous downward trend.
An increased response length is not necessary for better performance. Although we observed a rise in response length in CDAS and Curriculum Sampling, the response length in Dynamic Sampling stabilized after about 15 steps, yet the average accuracy on benchmarks continued to grow.
5.3 Ablations
Since CDAS is built on a fixed-point system, the choice of initial values can affect its convergence. We investigate this impact by removing the warmup phase and starting sampling from the first step.
Figure 7: Ablation study of the warm-up phase
As observed in fig. 7, the training curve exhibits significant fluctuations in the early stages of training (the first $20$ steps) without the warmup stage. Specifically, in the first $6$ steps, the removal of the warmup phase leads to a notable decrease in response length along with a rapid increase in reward, suggesting that the model has over-learned from simpler samples. However, as the convergence of the entire fixed-point system is guaranteed, the two training curves gradually overlap, further corroborating the stability of our framework.
Figure 8: Ablation study of symmetric sampling.
We further investigate the effectiveness of alignment-based symmetric sampling. For comparison, we directly select the $|B|$ problems with the smallest value of $A$ for the rollout phase.
As seen in fig. 8, after removing symmetric sampling, CDAS consistently outperforms the Random Sampling baseline. However, there is still an obvious decline in accuracy, especially in the later stages of training (after $100$ steps). We attribute this phenomenon to the accumulated estimation errors in model competence due to the imbalanced sampling. In the early stages of training, imbalanced sampling can cause the model’s competence to be either overestimated or underestimated. This discrepancy from the true competence accumulates as the number of training steps increases and hinders the convergence of CDAS, impairing the performance in the later stages.
6 Analysis and Discussion
In this section, we conduct a comprehensive analysis and discussion of CDAS. We explore the statistical properties of CDAS in terms of sample utility, investigate its advantages over pass rate-based methods, and validate its generalization to code generation tasks along with effectiveness on larger-scale models.
6.1 Utility of the Sampled Problems
Figure 9: The proportion of zero-gradient problems in the sampled batch.
From the perspective of the optimization objective of GRPO, the superior performance of Dynamic Sampling can be attributed to its filtering out of samples that do not contribute to model gradientsyu2025dapo (i.e., those with a pass rate of 00 or $1$). Although CDAS does not explicitly constrain the pass rate in problem selection, its alignment-based symmetric sampling inherently mitigates the issue of oversampling the zero-gradient problems to some extent. As illustrated in fig. 9, the proportion of such zero-gradient problems within the batches sampled by CDAS is consistently lower than that of the other baselines, proving that CDAS can effectively improve the utility of sampled problems. We also observe that the proportion of zero-gradient problems in CDAS exhibits a rapid decline during the early stages of training, followed by a slight increase in the later stages. The sharp decrease in the initial phase can be attributed to the swift correction of problem difficulty from its initial values. The modest rise in the later phase is mainly due to the increasing proportion of zero-gradient problems in the whole MATH training set, leading to more problems with a pass rate of 0 sampled in the batch $B^{+}$.
6.2 Problem Difficulty vs. Pass Rate
Since the problem difficulty in CDAS derived from the pass rates, we explore the relationship between them. As illustrated in fig. 10, problem difficulty and pass rate exhibit an overall negative correlation where problems with lower difficulty tend to have higher pass rates. Interestingly, we find that even among problems with the same pass rate, there can still be considerable differences in their estimated difficulties. To further investigate this phenomenon, we randomly selected two problems with a pass rate of 1 at the final sampling step.
Figure 10: Problem difficulty vs. pass rate in CDAS.
As shown in fig. 10, problem A required 25 samplings to reach a pass rate of 1, whereas problem B achieved a pass rate of 1 after only 6 samplings. Despite both having the same final pass rate in rollout, Problem A is noticeably more difficult than Problem B, as indicated by its average accuracy of $32$ inferences is $0.6875$, which is much lower than that of Problem B (Avg@32 = $1.0$). This validates that CDAS, by leveraging historical information, provides a more accurate and robust measure of problem difficulty.
6.3 Generalization
Extension to Code Generations To further investigate the effectiveness of our approach in diverse domains, we apply CDAS to GRPO training on code generation tasks. Specifically, We aggregate open-source data from Appshendrycks2021apps , Tacoli2023taco , and CodeContestsli2022codecontest , and perform GRPO training on Qwen2.5-Coder-7Bhui2024qwen2d5coder for 100 steps. Evaluation results on the LiveCodeBench v5 jain2025livecodebench are shown in LABEL:fig:lcb. Despite the challenges posed by this task for 7B size models, CDAS consistently outperforms the vanilla GRPO baseline in both pass@8 and pass@1 metrics after just 50 steps.
Effectiveness on Larger LLMs In addition to the strong performance observed on Qwen2.5-7B, we further validate the effectiveness of CDAS on larger LLMs. Specifically, we conduct training on Qwen2.5-14Byang2024qwen2p5 with a batch size of 256 for 200 steps, matching the computational budget of our main experiments. The average accuracy of CDAS and Random Sampling is reported in LABEL:fig:14b. We find that CDAS achieves substantial improvements over the Random Sampling baseline by $+1.47\%$, which is even greater than the improvement observed on the 7B model $+1.27\%$, showing effectiveness on larger models.
7 Conclusion
We present Competence-Difficulty Alignment Sampling (CDAS), a novel sampling strategy for RL training in LLM reasoning. CDAS addresses the limitations of existing methods by modeling problem difficulty as a trajectory of performance discrepancies to provide more stable estimations and explicitly aligning it with model competence at each training step throughout a fixed-point system. Extensive experiments on mathematical reasoning benchmarks demonstrate the superiority of CDAS in both accuracy and efficiency to powerful baselines. Our results highlight the importance of dynamically matching problem difficulty to model competence for efficient RL training.
Appendix A Training Details
The prompt template used for our zero RL training is shown below:
[…]
Table 2 presents the key configuration used for our Qwen2.5-7B experiment. Training was conducted on a single node with 8 A100 GPUs, and the model was trained for 110 steps using the veRL library [31]. For Qwen2.5-14B experiment, training was conducted on four nodes, each equipped with 8 A100 GPUs. The model was trained for 200 steps with a batch size of 256. Other configurations were kept consistent with those used in the Qwen2.5-7B experiment.
1python -m verl.trainer.main_ppo \
2 algorithm.adv_estimator=grpo \
3 data.train_batch_size=1024 \
4 data.val_batch_size=500 \
5 data.max_prompt_length=1024 \
6 data.max_response_length=4096 \
7 actor_rollout_ref.actor.optim.lr=5e-7 \
8 actor_rollout_ref.model.use_remove_padding=True \
9 actor_rollout_ref.actor.ppo_mini_batch_size=256 \
10 actor_rollout_ref.actor.ppo_micro_batch_size_per_gpu=4 \
11 actor_rollout_ref.actor.use_kl_loss=True \
12 actor_rollout_ref.actor.kl_loss_coef=0.001 \
13 actor_rollout_ref.actor.entropy_coeff=0.001 \
14 actor_rollout_ref.actor.clip_ratio=0.2 \
15 actor_rollout_ref.actor.kl_loss_type=low_var_kl \
16 actor_rollout_ref.model.enable_gradient_checkpointing=True \
17 actor_rollout_ref.actor.fsdp_config.param_offload=False \
18 actor_rollout_ref.actor.fsdp_config.grad_offload=False \
19 actor_rollout_ref.actor.fsdp_config.optimizer_offload=False \
20 actor_rollout_ref.rollout.temperature=1.0 \
21 actor_rollout_ref.rollout.log_prob_micro_batch_size=160 \
22 actor_rollout_ref.rollout.tensor_model_parallel_size=4 \
23 actor_rollout_ref.rollout.enable_chunked_prefill=False \
24 actor_rollout_ref.rollout.max_num_batched_tokens=5120 \
25 actor_rollout_ref.rollout.name=vllm \
26 actor_rollout_ref.rollout.gpu_memory_utilization=0.8 \
27 actor_rollout_ref.rollout.n=8 \
28 actor_rollout_ref.ref.log_prob_micro_batch_size=160 \
29 actor_rollout_ref.ref.fsdp_config.param_offload=True \
30 algorithm.kl_ctrl.kl_coef=0.001 \
31 critic.ppo_micro_batch_size_per_gpu=4 \
32 trainer.critic_warmup=0 \
33 trainer.n_gpus_per_node=8 \
34 trainer.nnodes=1 \
35 trainer.remove_clip=False
Table 2: Key configuration for our experiment.
Appendix B Evaluation Benchmarks
We evaluate mathematical problem-solving ability using a variety of well-known benchmarks, ranging from middle school to Olympiad-level difficulty:
  • AIME (American Invitational Mathematics Examination): A challenging high school level competition featuring 30 problems with integer answers. It is a well-established benchmark for evaluating advanced mathematical reasoning. Available at https://huggingface.co/datasets/math-ai/aime25 and https://huggingface.co/datasets/math-ai/aime24.
  • AMC (American Mathematics Competitions): A benchmark focuses on mathematical problem-solving skills for middle and high school students. Problems range in difficulty and emphasize logical reasoning. Dataset available at https://huggingface.co/datasets/math-ai/amc23.
  • MATH500 [12]: A curated subset of the larger MATH dataset, consisting of 500 problems that span various mathematical fields such as algebra, geometry, and number theory.
  • MinervaMath [15]: A benchmark used to evaluate the performance of large language models on detailed, multi-step quantitative reasoning problems.
  • OlympiadBench [16]: A benchmark specifically designed for assessing models on Olympiad-level mathematics, including problems from national and international math competitions.
  • GSM8K [17]: A dataset of grade school math word problems, commonly used to test arithmetic and step-by-step reasoning in models.
Appendix C Convergence Analysis
Recall the following coupled fixed-point system:
$\begin{cases}D^{*}(x)=\sigma(C^{*}-D^{*}(x))-S^{*}(x),\quad x\in\mathbb{X}\\ C^{*}=-\mathbb{E}_{x}[D^{*}(x)]\end{cases}$ (13)
where $\sigma(z)=\frac{1}{1+e^{-z}}$ is the sigmoid function, and $S^{*}(x)$ is a constant with $x$.
Let $\mathbf{D}=(D(x_{1}),D(x_{2}),...,D(x_{N}))^{\top}$ denote the vector of problem difficulties, and $C$ denote the model competence. The system can be written as a joint mapping $\mathcal{F}$ in $(N+1)$-dimensional space:
$\begin{cases}D_{n+1}(x)=\sigma(C_{n}-D_{n}(x))-S^{*}(x)\\ C_{n+1}=-\frac{1}{N}\sum_{i=1}^{N}D_{n+1}(x_{i})\end{cases}$
or equivalently, $(\mathbf{D}_{n+1},C_{n+1})=\mathcal{F}(\mathbf{D}_{n},C_{n})$.
The sigmoid function $\sigma(z)$ is Lipschitz continuous[36] with constant $L_{\sigma}=\frac{1}{4}$. For any $D(x),D^{\prime}(x)$ and $C,C^{\prime}$, we have:
$|\sigma(C-D(x))-\sigma(C^{\prime}-D^{\prime}(x))|\leq\frac{1}{4}\left(|C-C^{\prime}|+|D(x)-D^{\prime}(x)|\right)$
For two states $(\mathbf{D},C)$ and $(\mathbf{D}^{\prime},C^{\prime})$, define the distance as
$\|(\mathbf{D},C)-(\mathbf{D}^{\prime},C^{\prime})\|=\max\left\{\max_{x}|D(x)-D^{\prime}(x)|,|C-C^{\prime}|\right\}$
Then,
$|D_{n+1}(x)-D^{\prime}_{n+1}(x)|\leq\frac{1}{4}\left(|C_{n}-C^{\prime}_{n}|+|D_{n}(x)-D^{\prime}_{n}(x)|\right)$
Taking the maximum over all $x$, we obtain
$\max_{x}|D_{n+1}(x)-D^{\prime}_{n+1}(x)|\leq\frac{1}{4}\left(|C_{n}-C^{\prime}_{n}|+\max_{x}|D_{n}(x)-D^{\prime}_{n}(x)|\right)\leq\frac{1}{2}\delta_{n}$
where $\delta_{n}=\max\left\{\max_{x}|D_{n}(x)-D^{\prime}_{n}(x)|,|C_{n}-C^{\prime}_{n}|\right\}$.
For the update of $C$, we have
$|C_{n+1}-C^{\prime}_{n+1}|=\left|-\frac{1}{N}\sum_{x}D_{n+1}(x)+\frac{1}{N}\sum_{x}D^{\prime}_{n+1}(x)\right|\leq\max_{x}|D_{n+1}(x)-D^{\prime}_{n+1}(x)|$
Therefore, the joint mapping satisfies.
$\|(\mathbf{D}_{n+1},C_{n+1})-(\mathbf{D}^{\prime}_{n+1},C^{\prime}_{n+1})\|\leq\frac{1}{2}\|(\mathbf{D}_{n},C_{n})-(\mathbf{D}^{\prime}_{n},C^{\prime}_{n})\|$
That is, the contraction constant is $\frac{1}{2}<1$.
Since the joint mapping is a contraction in the $(N+1)$-dimensional space[37], by the Banach fixed-point theorem, the sequence $(\mathbf{D}_{n},C_{n})$ converges to a unique fixed point $(\mathbf{D}^{*},C^{*})$, regardless of the initial values.
Appendix D Full Experimental Results
LABEL:fig:7bfullresults displays the full results for the Qwen2.5-7B model, while LABEL:fig:14bfullresults displays the corresponding outcomes for the Qwen2.5-14B model.
Limitations
(a) Easy-biased dataset
(b) Hard-biased dataset
Figure 19: The impact of dataset difficulty distribution.
In our main experiments, the difficulty of the problems in MATH is distributed around the model’s capabilities. However, the difficulty distribution of the dataset may be biased in practical scenes. To explore the impact of the overall difficulty of the dataset on the training of CDAS, we sample an extremely easy subset containing 3000 problems with the highest pass rate and an extremely difficult subset containing 3000 problems with the lowest pass rate based on the results of Qwen2.5-7B. We compare CDAS with Prioritized Sampling and Random Sampling, training them for 30 steps.
We observe an interesting phenomenon that when the overall difficulty of the dataset is extremely biased, there is no obvious difference between different sampling strategies (see fig. 19). We conjecture that one of the prerequisite for sampling strategies to be effective in the RL training is that the difficulty distribution of the dataset does not deviate too much from the model’s capabilities.